javonarnold:

The angles below are supplementary. What is the value of x? (1 point) A pair of supplementary angles is shown. One angle measures 6x plus 48 degrees, and the other angle measures 60 degrees. 2 7 12 18

Devum:

Supplementary= adding up to 180 degrees 6x+48+60=180 6x+108=180 -108 -108 6x=72 6x/6 = 72/6 x=12

javonarnold:

thanks

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold thanks \(\color{#0cbb34}{\text{End of Quote}}\) Happy to help! :)

Devum:

get some sleep tho lol its 12:30 where i am and u shouldnt be doing work at that kinda time

javonarnold:

i kno but me i always wanna stay ahead of my work this semesiter for i wont fail

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold i kno but me i always wanna stay ahead of my work this semesiter for i wont fail \(\color{#0cbb34}{\text{End of Quote}}\) yeah ik me too but like u can do this in the morning lol

javonarnold:

true:)

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold true:) \(\color{#0cbb34}{\text{End of Quote}}\) Yess so gts now dude 😂

javonarnold:

but if i need help with anything else u mind if i hit u up

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold but if i need help with anything else u mind if i hit u up \(\color{#0cbb34}{\text{End of Quote}}\) Nah go right on ahead hit me up at @Devum or @Convert which is my other account and ill try to get back to u asap

javonarnold:

im not goin to sleep until i get these three questions done

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold im not goin to sleep until i get these three questions done \(\color{#0cbb34}{\text{End of Quote}}\) Alright well if you need help with anyone jus let me know lol

javonarnold:

u mind helping me with these three questions tho u just be confused sometimes and get stress out easily

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold u mind helping me with these three questions tho u just be confused sometimes and get stress out easily \(\color{#0cbb34}{\text{End of Quote}}\) Yeah ofc dude jus post em

javonarnold:

ok

javonarnold:

A running track in the shape of an oval is shown. The ends of the track form semicircles. A running track is shown. The left and right edges of the track are identical semicircles. The top and bottom edges of the track are straight lines. The track has width 42 m and length of one straight edge 110 m. What is the perimeter of the inside of the track? (Ï€ = 3.14) (1 point) 304.00 m 351.88 m 912.37 m 1,604.74 m

javonarnold:

if u need the picture instead ill try to give it to u

Devum:

Yeah the picture wouldn help lol

javonarnold:

k

Devum:

Thx dude I appreciate it

javonarnold:

np but im trying to put the sc on here but its hard

javonarnold:

sorry if im taking long

javonarnold:

Devum:

Ohh okay I see so 1st you find the perimeter of a rectangle so 2L*2W L being length and W being width

Devum:

Then the circumference is 2*the diameter then multiply that times pi

javonarnold:

ok give me a sec

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold ok give me a sec \(\color{#0cbb34}{\text{End of Quote}}\) Alright

Devum:

And the 2 halfs on the end make a full circle btw lol

javonarnold:

ok im confused and gettting stress out and i kno im taking up some much of your time so u dont gotta help me

Devum:

Nooo I'll help you

Devum:

Here I'll give you a step by step answer hold up though I'm doing my own hw but I'm almost done

javonarnold:

ok

Devum:

Okay so now you see we have 42 and 110 as our sides: 2*42+2*110=304 Now we have the perimeter of the square, but we need to subtract the 2, 42 sides because it does not border the side. 304-2*42=220 Now that we have that we need to find the circumference (perimeter) of a circle: The formula is diameter*pi: 42*3.14=131.88 Now that we have to circumference, even though it is a semi-circle, because we have 2 halves they equal a whole. Finally we have our parts and we add them up. 220+131.88=351.88 So from this we can determine it'd be the 2nd answer. 351.88.

javonarnold:

@Devum u still here

Devum:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @javonarnold @Devum u still here \(\color{#0cbb34}{\text{End of Quote}}\) Yeah I'm still here. What's up?

Devum:

I was boutta head to sleep tho its gettin late ;-;

javonarnold:

ok